# Vejde: A Framework for Inductive Deep Reinforcement Learning Based on Factor Graph Color Refinement

### v1

@@ 1 Introduction
We are interested in two topics: Deep reinforcement learning (RL) for problem domains that fit relational data models, and agents that can generalize to classes of problems. This interest mainly stems from our experiences in researching automated network incident response, where both of the aforementioned qualities are important for practical use, though not always prioritized (Wolk et al., 2022; Nyberg & Johnson, 2024; Thompson et al., 2024). The two topics are interleaved, in that incorporating structure to the traditional Markov decision process (MDP) formalism typically used in RL is in itself a method for improving agent generalization (Van Otterlo, 2009; Mohan et al., 2024; Kirk et al., 2023), though neither topic necessitates the other. To this end, we have developed a small reinforcement learning library which we have named Vejde. [^1] Vejde is the Swedish name of the flower Isatis tinctoria, traditionally refined to produce indigo dye for textiles. Vejde combines graph neural networks and reinforcement learning to incorporate data structures which can be observed or defined in different problem areas, and to produce neural policy functions that can be used for classes of problems defined with a shared data description language.
@@
A decision agent policy is a function which takes the state of a decision problem, usually modeled as a MDP or partially observable MDP, as input and produces a probability distribution over actions that are possible in the given state. An optimal policy will produce actions that maximize the expected reward; the assigned metric of success at the task the MDP models. Our aim is to find inductive policies that can perform well across a domain, or class, of problems. This has long been a goal within the realm of relational RL (Dzeroski et al., 2001; Van Otterlo, 2009), which combines reinforcement learning with elements of symbolic logic, but ideas for improving RL generalization have also come from more recent sources in deep learning research (Kirk et al., 2023).
@@
Relational database systems and data modeling remains a popular option for data storage in several areas, [^2] According to https://db-engines.com/en/ranking. which motivates investigating machine learning methods that can utilize the structure relational databases provide. Relational databases use SQL (Chamberlin & Boyce, 1974) and to varying degrees conform with a formal relational model (Codd, 1970). Previous work has investigated methods for applying supervised learning to relational databases (Fey et al., 2024b) using graph neural networks. From that perspective, this work concerns instead using RL and GNNs with relational databases. Another way of regarding a relational database is as a set of facts about entities described using a typed first-order logical language, where every possible set of facts given the language is a discrete state of some problem domain.
@@
One way of describing a policy for such a logical state representation is as a decision list. Each entry in the decision list defines a set of logical conditions, defined in the same language or data model as the database, that when fulfilled leads to an agent choosing an action (Fern et al., 2006). There exists approaches, both symbolic (Fern et al., 2006) and neuro-symbolic (Hazra & Raedt, 2023) which aim to find and represent the decision list in an explicit and human-readable form. We instead opt to use message-passing neural networks to implicitly encode the rules of a decision list. This choice trades interpretability for flexibility and ease of numerical optimization, as is typical with deep learning methods. The abilities, and limitations, of MPNNs to express logical classifiers has been covered by Barceló et al. (2020) and Grohe (2021).
@@
Our work builds upon previous work in the area of symbolic, and neuro-symbolic, decision learning, both within the realm of automated planning through machine learning (Ståhlberg et al., 2023; Chen & Thiébaux, 2024) and deep relational reinforcement learning (Janisch, 2024; Sharma et al., 2023). Unlike problems common in symbolic planning research, we focus on problems with uncertain outcomes, formalized as MDPs, and which we assume may include continuous variables.
@@ 2 Background
This section covers topics related to knowledge representation, graph neural networks and reinforcement learning that are relevant to this work. We should note that the focus of this work is not on how to design, or learn, data abstractions for a problem domain. We recognize the importance of this work in creating agents, but assume that for a given problem there exists a data model of object classes, attributes and relations.
@@ 2.1 Knowledge Representation
Predicate logic, or first-order logic, has long been used to describe situations and dynamics of decision problems (Reiter, 2001; Russell & Norvig, 2010). Through this text, we will use components of first-order logic to describe the states of problems. We thus use the following definitions: Predicates are used to represent properties or relations. We will denote predicate symbols with capital letters such as $P$ or $Q$. An atom is a predicate symbol combined with tuple of variables, denoted with lowercase letters. Predicate symbols have an arity which defines the length of tuples it can be applied to. We will use the term object when referring to instantiations of variables, denoted with lowercase letters and a subscript, like $x_{1}$ or $y_{1}$. A fact refers to a ground atom where all variables have been substituted by objects, as in $P(x_{1})$ or $Q(x_{1},y_{1})$. Function symbols map objects or tuples of objects to other elements in the language. Given a function symbol $Z$, we could for instance make the statement $Z(x_{1})=30$. Lastly, we assume a typed logic. This means that each predicate is also associated with a tuple that specifies the types of the variables. Each object thus has a type or class. Any ground atom that is the result of combining a predicate symbol with a tuple of objects with mismatched types is defined as false. We will simply use the letter used to identify the object to indicate its type, meaning that $x_{1}$ and $y_{1}$ can be read as belonging to different object classes. All the definitions form a language, which can be specific to a particular problem domain.
@@ 2.2 Markov Decision Processes
We choose to represent decision problems as Markov decision processs (Puterman, 1994). A MDP is a formalization of a system where an agent makes sequential decisions, under the assumption of potential uncertainty about the decision outcomes. A reward or cost value defines a measure of success at the task the agent should perform. MDPs are assumed to have the Markov property that the state and reward dynamics are dependent only on the immediate previous state and action. The agent is assumed to follow a policy, $\pi$, by which it selects actions given the state it observes. For a finite MDP, an optimal policy maximizes the expected discounted return, $G_{t}=\sum^{T}_{i=t}\gamma^{i}r_{i}$. The term relational MDP has been used by several authors to describe MDPs where the problem state is structured, factored or explicitly represented using symbolic logic (Guestrin et al., 2003; Diuk et al., 2008; Dzeroski et al., 2001; Van Otterlo, 2009; Kersting & Driessens, 2008). A relational MDP can be described as collection of problems, each individually a MDPs, which all share a common description language.
@@ 2.3 Reinforcement Learning
$L_{p}(s_{t},a_{t},\theta)=\min\left(r(\theta)\cdot A(s_{t},a_{t}),\;\;\text{clip}\left(r(\theta),1-\epsilon,1+\epsilon\right)\cdot A(s_{t},a_{t})\right)$
where $r(s_{t},a_{t},\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi^{\prime}_{\theta}(a_{t}|s_{t})}$, a ratio between the old and updated policy output which is used to clip the loss. $A$ is the action advantage, which may be calculated in different ways, but is typically defined as the difference between the state-value and action-value for the action.
@@ 3 Implementation
Vejde is formed by four primary components: First is representing data in the problem state using predicate logic, second is the representation of the data base as a bipartite graph, third is the graph neural network and lastly a prediction head which uses the object “colors” to produce action probabilities and a state value estimation. A key implementation detail is that the size of both the input and the number of actions can change between steps.
@@ 3.1 States & Actions
We represent the state of the MDP in a given time step as a set of facts expressed in the domain language. A particular state could thus be described as the set
$\left\{P(x_{1}),Q(x_{1},y_{1}),Z(y_{1})=5,Q(x_{2},y_{1}),C\right\}$
given the predicates $\left\{P,Q,Z,C\right\}$ and objects $\left\{x_{1},y_{1},y_{2}\right\}$. [^6] Since the truthiness and values of literals are assumed to be variable, it would be more correct to denote each fact with a time or step. We do not, however, since our calculations only ever involve the current state. As a delimitation, function symbols are assumed to always map objects to singular values in $\mathbb{R}$. Facts that are explicitly known to be false are not included in the state data base. If we maintain the assumption of full observability in the environment, there is no information lost by making this choice, as we do not need to distinguish between a fact being negative or missing from the data base (Reiter, 1977). This choice makes the state description significantly smaller than if negative literals are included. Actions in the MDP are represented as ground atoms from the same language as the state by assigning a subset of predicates in the domain language as action symbols, $A\subset P$. As an example, given the action symbols $\left\{A_{1},A_{2}\right\}$, which can have different arities and type restrictions, a possible set of actions in the above state may be $\left\{A_{1}(x_{1}),A_{1}(x_{2}),A_{2}(x_{1})\right\}$.
@@
3.1.1 State Graph
In order to generate object representations through graph color refinement, we first represent the state database as a biparitite graph, which consists of two distinct sets of nodes, $V$ and $U$. $V$ is a set of all facts in the state and $U$ is a set of observed objects. Thus, for the state shown in Section 3.1, $U=\left\{x_{1},y_{1},x_{2}\right\}$ and $V=\left\{P(x_{1}),\ Q(x_{1},y_{1}),\ Z(y_{1})=5,\ Q(x_{2},y_{1}),\ C\right\}$. A set of edges $E$ relate facts in $V$ to objects in $U$. Each edge is formed by an object being present in the object tuple of a fact. Following the representation in Chen et al. (2024), each edge between an object $u$ and a literal $v$ is associated with the position of $u$ in the tuple of objects of $v$. A rendering of the state graph representation is shown in 1(a), along with possible actions in 1(b).
@@ 3.2 Vector Encoding
All elements of the state are encoded into the vector space $\mathbb{R}^{D}$ using embedding vectors. The encoded elements form two multisets of node embeddings $K=\left\{\left\{k_{v}:v\in V\right\}\right\}$ and $H=\left\{\left\{h_{u}:u\in U\right\}\right\}$, corresponding to the factor nodes in $V$ and object nodes in $U$. These form the initial set of node colors.
@@ 3.2.1 Literals
We include two types of ground expressions in our data model; ground atoms, e.g. $P(x_{1}),Q(x_{1},y_{1})$, and grounded function symbols that map objects to a real value e.g. $Z(y_{1})=1$. Ground atoms are encoded by the embedding vector of the predicate symbol. Function atoms are encoded as a linear scaling of the function symbol’s embedding vector by the value it is equal to.
We can represent this in a single expression by defining the function $\nu:V\rightarrow\mathbb{R}$ that maps all elements of $V$ to a value from $\mathbb{R}$. If the element is a ground atom, the value is defined to be 1, and it is an expression with a function symbol it returns the value of the function. Thus, we can define the embedding vector $k_{v}$ for a literal $v$ as
$k_{v}=\nu(v)\cdot E_{P}(f_{V\rightarrow{}P}(v)),\ E_{P}:P\rightarrow\mathbb{R}^{D}.$
Expressions with nullary atoms are treated with an additional step. We denote the set of these literals as $G\subseteq V$, and assume that nullary literals influence all objects. The encoding vectors of the nullary literals are aggregated and added to the initial embedding vector of each object. We use the function
@@ 3.2.2 Objects
To avoid using object identifiers as features, objects in the database are mapped to their respective types through a function $f_{U\rightarrow{}T}:U\rightarrow T$, where $T$ is the set of types defined in the language. The type identifiers are then mapped to vectors in $\mathbb{R}^{D}$. The embedding vector $h_{u}$ for an object $u$ is thus defined as
$h_{u}=E_{T}(f_{U\rightarrow{}T}(u)),\ E_{T}:T\rightarrow\mathbb{R}^{D}.$
@@ 3.2.3 Positions
We encode the positions of terms in literals by mapping argument positions to vectors in $\mathbb{R}^{D}$. The function $f_{U,V\rightarrow{}N}$ takes a fact and an object, and maps them to the set of arities for predicates in the language, $N$.
$e_{uv}=E_{N}(f_{U,V\rightarrow{}N}(u,v)),\ E_{N}:N\rightarrow\mathbb{R}^{D}.$
@@ 3.3 Message Passing
The message-passing scheme we use for color refinement is similar to previous work on neural message passing on factor graphs (Satorras & Welling, 2021). A single iteration of message passing consists of nodes in $K$ sending messages to neighboring nodes in $H$. $H$ is updated and sends messages back to $K$ which is then also updated. Each step of message passing uses a unique set of four neural networks, which we denote with $\phi$. Incoming messages to a node are aggregated by taking the element-wise maximum over the node neighborhood, $\max_{x\in N(y)}:D\rightarrow D$. While a $\max$ aggregation is theoretically less expressive than a sum aggregation (Xu et al., 2019), it has been found to work well in practice in previous work (Janisch, 2024; Ståhlberg et al., 2023). Messages from $U$ to $V$, and the reverse, are defined as
$m^{(i+1)}_{v\rightarrow u}=\phi^{(i)}_{V\rightarrow U}(h^{(i)}_{u}\mathbin{\|}k^{(i)}_{v})$
$m^{(i+1)}_{u\rightarrow v}=\phi^{(i)}_{U\rightarrow V}(k^{(i)}_{v}\mathbin{\|}h^{(i+1)}_{u}\mathbin{\|}e_{uv})$
and representation vectors for object nodes in $U$ as
@@ 3.4 Policy
A policy function can in a more general view be regarded as classifier, which takes a state representation and assigns a probability to each action that is possible given the state. In this setting, the state is represented as a set of facts, which are encoded using embedding vectors and message passing to a latent state $H$ of object embeddings. Thus, from the multiset $H$, the policy function should assign a probability to each action that is possible given the objects in $U$ and the action symbols $A$, as described in Section 3.1. Since we limit action symbols to be nullary or unary, all actions contain a predicate symbol, $a\in A\subset P$, and a single object, $u\in U$. Nullary actions have no arguments by definition, but for ease of implementation we define a null object, $\varnothing$, which is always selected for nullary actions. We denote the joint action probability as $\pi(a,u|H)$, such that $\pi(a_{1},u_{1}|H)$ would be the probability of the action $a_{1}(u_{1})$ given the set of object embeddings, $H$. The full joint action probability can be predicated as is, or factorized in three different ways, which lead to different agent properties (Nyberg & Johnson, 2024). [^7] The main property that changes is the dependency on global readouts. Sampling an object first and then an action symbol based on that choice can be done without considering the entire graph. For this work, we solely evaluate the factorization $\pi(a,u|H)=\pi_{A}(a|H)\cdot\pi_{U}(u|a,H)$, which corresponds to first sampling a predicate symbol, followed by an object. For notation brevity, we will omit the conditional of $H$ from here onward. We calculate the action symbol probability, $\pi_{A}(a)$, as a weighted sum of conditional action probabilities over objects
@@
where
$\pi_{A}(a|u):=\text{Softmax}^{A}({W\cdot h_{u},\ a}),\ W\in\mathbb{R}^{|A|\times D}$
and
$\pi_{U}(u):=\text{Softmax}^{U}(W\cdot H,u),\ W\in\mathbb{R}^{D}.$
The probability of selecting an object given an action symbol, $\pi_{U}(u|a)$, is calculated as
$\pi_{U}(u|a):=\text{Softmax}^{U}(H\cdot W_{a},u),\ W\in\mathbb{R}^{D\times|A|}$
where $W_{a}$ is the column vector in $W$ identified by $a$. We use $\text{Softmax}^{D}(x,i)=\nicefrac{{e^{x_{i}}}}{{\sum^{D}_{j}e^{x_{j}}}}$ to denote the $i$-th element of a vector $x$ where the softmax function is applied over dimension $D$.

### v2

@@ 1 Introduction
We are interested in two topics: Deep reinforcement learning (RL) for problem domains that fit relational data models, and agents that can generalize to classes of problems. This interest mainly stems from our experiences in researching automated network incident response, where both of the aforementioned qualities are important for practical use, though not always prioritized (wolk2022cage; DBLP:conf/csr2/NybergJ24; DBLP:conf/autonomouscyber/ThompsonCHM24). The two topics are interleaved, in that incorporating structure to the traditional Markov decision process (MDP) formalism typically used in RL is in itself a method for improving agent generalization (vanotterloLogicAdaptiveBehavior2009a; 10.1613/jair.1.15703; DBLP:journals/jair/KirkZGR23), though neither topic necessitates the other. To this end, we have developed a small reinforcement learning library which we have named Vejde. [^1] Vejde is the Swedish name of the flower Isatis tinctoria, traditionally refined to produce indigo dye for textiles. Vejde combines graph neural networks and reinforcement learning to incorporate data structures which can be observed or defined in different problem areas, and to produce neural policy functions that can be used for classes of problems defined with a shared data description language.
@@
A decision agent policy is a function which takes the state of a decision problem, usually modeled as a MDP or partially observable MDP, as input and produces a probability distribution over actions that are possible in the given state. An optimal policy will produce actions that maximize the expected reward; the assigned metric of success at the task the MDP models. Our aim is to find inductive policies that can perform well across a domain, or class, of problems. This has long been a goal within the realm of relational RL (DBLP:journals/ml/DzeroskiRD01; vanotterloLogicAdaptiveBehavior2009a), which combines reinforcement learning with elements of symbolic logic, but ideas for improving RL generalization have also come from more recent sources in deep learning research (DBLP:journals/jair/KirkZGR23).
@@
Relational database systems and data modeling remains a popular option for data storage in several areas, [^2] According to https://db-engines.com/en/ranking. which motivates investigating machine learning methods that can utilize the structure relational databases provide. Relational databases use SQL (10.1145/800296.811515) and to varying degrees conform with a formal relational model (10.1145/362384.362685). Previous work has investigated methods for applying supervised learning to relational databases (fey_position_2024) using graph neural networks. From that perspective, this work concerns instead using RL and GNNs with relational databases. Another way of regarding a relational database is as a set of facts about entities described using a typed first-order logical language, where every possible set of facts given the language is a discrete state of some problem domain.
@@
One way of describing a policy for such a logical state representation is as a decision list. Each entry in the decision list defines a set of logical conditions, defined in the same language or data model as the database, that when fulfilled leads to an agent choosing an action (DBLP:journals/jair/FernYG06). There exists approaches, both symbolic (DBLP:journals/jair/FernYG06) and neuro-symbolic (DBLP:conf/pkdd/HazraR23) which aim to find and represent the decision list in an explicit and human-readable form. We instead opt to use message-passing neural networks to implicitly encode the rules of a decision list. This choice trades interpretability for flexibility and ease of numerical optimization, as is typical with deep learning methods. The abilities, and limitations, of MPNNs to express logical classifiers has been covered by Barceló et al. (2020) and Grohe (2021).
@@
Our work builds upon previous work in the area of symbolic, and neuro-symbolic, decision learning, both within the realm of automated planning through machine learning (DBLP:conf/kr/StahlbergBG23; DBLP:conf/nips/ChenT24) and deep relational reinforcement learning (janisch2024applications; DBLP:conf/uai/SharmaAMS23). Unlike problems common in symbolic planning research, we focus on problems with uncertain outcomes, formalized as MDPs, and which we assume may include continuous variables.
@@ 2 Background
This section covers topics related to knowledge representation, graph neural networks and reinforcement learning relevant to this work. We emphasize that our focus is not on how to design or learn data abstractions for a problem domain, and assume that for a given problem there already exists a data model of object classes, attributes and relations.
@@ 2.1 Knowledge Representation
Predicate logic, or first-order logic, has long been used to describe situations and dynamics of decision problems (reiterKnowledgeActionLogical2001a; russel2010). Through this text, we will use components of first-order logic to describe the states of problems. We thus use the following definitions: Predicates are used to represent properties or relations. We will denote predicate symbols with capital letters such as $P$ or $Q$. An atom is a predicate symbol combined with tuple of variables, denoted with lowercase letters. Predicate symbols have an arity which defines the length of tuples it can be applied to. We will use the term object when referring to instantiations of variables, denoted with lowercase letters and a subscript, like $x_{1}$ or $y_{1}$. A fact refers to a ground atom where all variables have been substituted by objects, as in $P(x_{1})$ or $Q(x_{1},y_{1})$. Function symbols map objects or tuples of objects to other elements in the language. Given a function symbol $Z$, we could for instance make the statement $Z(x_{1})=30$. Lastly, we assume a typed logic. This means that each predicate is also associated with a tuple that specifies the types of the variables. Each object thus has a type or class. Any ground atom that is the result of combining a predicate symbol with a tuple of objects with mismatched types is defined as false. We will simply use the letter used to identify the object to indicate its type, meaning that $x_{1}$ and $y_{1}$ can be read as belonging to different object classes. All the definitions form a language, which can be specific to a particular problem domain.
@@ 2.2 Markov Decision Processes
We choose to represent decision problems as Markov decision processs (DBLP:books/wi/Puterman94). A MDP is a formalization of a system where an agent makes sequential decisions, under the assumption of potential uncertainty about the decision outcomes. A reward or cost value defines a measure of success at the task the agent should perform. MDPs are assumed to have the Markov property that the state and reward dynamics are dependent only on the immediate previous state and action. The agent is assumed to follow a policy, $\pi$, by which it selects actions given the state it observes. For a finite MDP, an optimal policy maximizes the expected discounted return, $G_{t}=\sum^{T}_{i=t}\gamma^{i}r_{i}$. The term relational MDP has been used by several authors to describe MDPs where the problem state is structured, factored or explicitly represented using symbolic logic (DBLP:conf/ijcai/GuestrinKGK03; DBLP:conf/icml/DiukCL08; DBLP:journals/ml/DzeroskiRD01; vanotterloLogicAdaptiveBehavior2009a; DBLP:conf/icml/KerstingD08). A relational MDP can be described as collection of problems, each individually a MDPs, which all share a common description language.
@@ 3 Implementation
Vejde is formed by four primary components described in this section: Representing the problem state using predicate logic, representing the data base as a bipartite graph, encoding the graph using a graph neural network and lastly using the object encoding vectors to produce action probabilities and a state value estimation. A key implementation detail is that the sizes of both the agent input and output is not fixed, and dependent on the number of facts and entities in the state.
@@ 3.1 States & Actions
We represent the state of the MDP in a given time step as a set of facts expressed in the domain language [^6] Since the truthiness and values of literals are assumed to be variable, it would be more correct to denote each fact with a time or step. We do not, however, since our calculations only ever involve the current state.. To reduce the size of the state data base, facts that are explicitly known to be false are not included. No information is lost through this choice if we maintain the assumption of full observability in the environment, as there is no need to distinguish between a fact being false or missing from the data base (DBLP:conf/adbt/Reiter77). Actions in the MDP are represented as ground atoms from the same language as the state by assigning a subset of predicates in the domain language as action symbols, $A\subset P$. An example state with actions is shown in Appendix A.
@@
3.1.1 Bipartite State Graph
In order to generate object representations through graph color refinement, we first represent the state database as a biparitite graph, which consists of two distinct sets of nodes, $V$ and $U$. $V$ is a set of all facts in the state and $U$ is a set of observed objects. A set of edges $E$ relate facts in $V$ to objects in $U$. Each edge is formed by an object being present in the object tuple of a fact. Following the representation in DBLP:conf/icaps/ChenTT24, each edge between an object $u$ and a literal $v$ is associated with the position of $u$ in the tuple of objects of $v$. A rendering of the state graph representation is shown in 1(a), along with possible actions in 1(b).
@@ 3.2 Vector Encoding
All elements of the state are encoded into the vector space $\mathbb{R}^{D}$ using embedding vectors. The encoded elements form two multisets of initial node embeddings $K=\left\{\left\{k_{v}:v\in V\right\}\right\}$ and $H=\left\{\left\{h_{u}:u\in U\right\}\right\}$, corresponding to the factor nodes in $V$ and object nodes in $U$. We also encode the positions of objects in the argument list of facts by mapping the positions to vectors in $\mathbb{R}^{D}$.
@@ 3.2.1 Literals
Two types of ground expressions are included in our data model; ground atoms, e.g. $P(x_{1}),Q(x_{1},y_{1})$, and grounded function symbols that map objects to a real value e.g. $Z(y_{1})=1$. Ground atoms are encoded by the embedding vector of the predicate symbol. Function atoms are encoded as a linear scaling of the function symbol’s embedding vector by the value it is equal to [^7] As a simplifying delimitation, function symbols are assumed to always map objects to singular values in $\mathbb{R}$.. We can represent the encoding of both types in a single expression by defining the function $\nu:V\rightarrow\mathbb{R}$ that maps all elements of $V$ to a value from $\mathbb{R}$. If the element is a ground atom, the value is defined to be 1, and it is an expression with a function symbol it returns the value of the function. Thus, we define the embedding vector $k_{v}$ for a literal $v$ as $k_{v}=\nu(v)\cdot E_{P}(f_{V\rightarrow{}P}(v)),\ E_{P}:P\rightarrow\mathbb{R}^{D}.$
Expressions with nullary atoms are handled with an additional step. We denote the set of these literals as $G\subseteq V$, and assume that nullary literals influence all objects. The encoding vectors of the nullary literals are aggregated and added to the initial embedding vector of each object. We use the function
@@ 3.2.2 Objects
To avoid using object identifiers as features, objects are mapped to their respective types through a function $f_{U\rightarrow{}T}:U\rightarrow T$, where $T$ is the set of types defined in the language. The type identifiers are then mapped to vectors in $\mathbb{R}^{D}$. The embedding vector $h_{u}$ for an object $u$ is thus defined as $h_{u}=E_{T}(f_{U\rightarrow{}T}(u)),\ E_{T}:T\rightarrow\mathbb{R}^{D}.$
@@ 3.2.3 Positions
Edge embeddings are defined as $e_{uv}=E_{N}(f_{U,V\rightarrow{}N}(u,v)),\ E_{N}:N\rightarrow\mathbb{R}^{D}$. The function $f_{U,V\rightarrow{}N}$ takes a fact and an object, and maps them to the set of arities for predicates in the language, $N$.
@@ 3.3 Message Passing
The message-passing scheme we use for color refinement is similar to previous work on neural message passing on factor graphs (Satorras and Welling, 2021). A single iteration of message passing consists of nodes in $K$ sending messages to neighboring nodes in $H$. $H$ is updated and sends messages back to $K$ which is then also updated. Each step of message passing uses a unique set of four neural networks, which we denote with $\phi$. Incoming messages to a node are aggregated by taking the element-wise maximum over the node neighborhood, $\max_{x\in N(y)}:D\rightarrow D$. While a $\max$ aggregation is theoretically less expressive than a sum aggregation (Xu et al., 2019), it has been found to work well in practice in previous work (janisch2024applications; DBLP:conf/kr/StahlbergBG23). Messages from $U$ to $V$ are defined as $m^{(i+1)}_{v\rightarrow u}=\phi^{(i)}_{V\rightarrow U}(h^{(i)}_{u}\mathbin{\|}k^{(i)}_{v})$ , and the reverse as $m^{(i+1)}_{u\rightarrow v}=\phi^{(i)}_{U\rightarrow V}(k^{(i)}_{v}\mathbin{\|}h^{(i+1)}_{u}\mathbin{\|}e_{uv})$ . Representation vectors for object nodes in $U$ are calculated at each iteration as
@@ 3.4 Policy
A policy function can in a more general view be regarded as classifier, which takes a state representation and assigns a probability to each action that is possible given the state. In this setting, the state is represented as a set of facts, which are encoded using embedding vectors and message passing to a latent state $H$ of object embeddings. Thus, from the multiset $H$, the policy function should assign a probability to each action that is possible given the objects in $U$ and the action symbols $A$, as described in Section 3.1. Since we limit action symbols to be nullary or unary, all actions contain a predicate symbol, $a\in A\subset P$, and a single object, $u\in U$. Nullary actions have no arguments by definition, but for ease of implementation we define a null object, $\varnothing$, which is the only selectable object for nullary actions. We denote the joint probability over action symbols and entities as $\pi(a,u|H)$, such that $\pi(a_{1},u_{1}|H)$ would be the probability of the action $a_{1}(u_{1})$ given the set of object embeddings, $H$. The joint action probability can be predicated in full, or factorized in three different ways, which lead to different agent properties (DBLP:conf/csr2/NybergJ24). [^8] The main property that changes is the dependency on global readouts. Sampling an object first and then an action symbol based on that choice can be done without considering the entire graph. For this work, we solely evaluate the factorization $\pi(a,u|H)=\pi_{A}(a|H)\cdot\pi_{U}(u|a,H)$, which corresponds to first sampling a predicate symbol, followed by an object. The conditional of $H$ will be omitted from here onward, for notation brevity. We calculate the action symbol probability, $\pi_{A}(a)$, as a weighted sum of conditional action probabilities over objects
@@
where $\pi_{A}(a|u):=\text{Softmax}^{A}({W\cdot h_{u},\ a}),\ W\in\mathbb{R}^{|A|\times D}$ and $\pi_{U}(u):=\text{Softmax}^{U}(W\cdot H,u),\ W\in\mathbb{R}^{D}.$
The probability of selecting an object given an action symbol, $\pi_{U}(u|a)$, is calculated as $\pi_{U}(u|a):=\text{Softmax}^{U}(H\cdot W_{a},u),\ W\in\mathbb{R}^{D\times|A|}$ where $W_{a}$ is the column vector in $W$ identified by $a$. We use $\text{Softmax}^{D}(x,i)=\nicefrac{{e^{x_{i}}}}{{\sum^{D}_{j}e^{x_{j}}}}$ to denote the $i$-th element of a vector $x$ where the softmax function is applied over dimension $D$.